### v2

# Simple and Provable Scaling Laws for the Test-Time Compute of Large Language Models

Abstract
We propose two simple yet principled algorithms that enjoy provable scaling laws for the test-time compute of large language models (LLMs), which require a black-box LLM and nothing else (e.g., no external verifier or reward model) for a minimalistic implementation.
  • The first one is a two-stage knockout-style algorithm: given an input problem, it first generates multiple candidate solutions, and then aggregate them for a final output, via a knockout tournament where pairwise comparisons among the candidates are conducted. Assuming that the LLM can generate a correct solution with non-zero probability and do better than a random guess in comparing a pair of correct and incorrect solutions, we prove theoretically that the failure probability of this algorithm decays to zero exponentially or by a power law (depending on the specific way of scaling) as its test-time compute grows.
  • The second one is a two-stage league-style algorithm, where each candidate solution is evaluated by its average win rate against multiple opponents, rather than eliminated upon loss to a single opponent. Under certain technical assumptions that are analogous to but more robust than those required by the knockout-style algorithm, we prove theoretically that the failure probability of the league-style algorithm also decays to zero exponentially as its test-time compute grows.
Through extensive experiments with two challenging benchmarks, namely GPQA and MMLU-Pro, we validate the proposed theories and demonstrate the outstanding scaling properties of both algorithms.
11footnotetext: Equal contributions.
1 Introduction
Despite the astonishing advancements of large language models (LLMs) in the past few years, they still face challenges with reliability and stability. This hinders their applications in high-stakes scenarios where a problem need to be solved with success probability $99.9\%$ rather than $90\%$. Similarly, in an LLM-based agentic workflow that involves solving many sub-problems, each of them need to be solved with high success probability in order to ensure the success of the complete workflow, since a single error in the process can lead to an incorrect final output. In these and many other similar scenarios, one is willing to boost the success probability by spending more test-time compute on LLM inference. Common approaches include letting the LLM generate a long chain-of-thought trajectory [38, 17, 26, 47], or asking the LLM to iteratively self-verify or self-refine the solutions that it has generated [7, 24, 8, 44, 45]. Another category of approaches is about sampling multiple candidate solutions and then choosing the best one via majority voting [3, 36, 2, 20], ranking with pairwise comparisons [12], or using an external verifier [14, 32] or trained reward model [5, 21, 49].
The primary goal of this work is to provide some theoretical insights into the full potential of scaling up the test-time compute for LLM inference [31, 42, 48, 39, 47, 11]. Throughout this work, we consider a generic problem formulation where an LLM-based algorithm is given an input problem and asked to output a solution. For conceptual simplicity, we evaluate any solution with a binary metric indicating whether it is correct or incorrect [^1] For example, a solution is “correct” if it contains the right reasoning process and final answer to a math problem, or it meets a set of criteria for an open-ended writing task. It is possible to expand the considered scenarios by adopting more general (discrete or continuous) metrics for evaluating a solution, though this is beyond the scope of our current focus.. In this setting, we seek to design practical and principled algorithms that enjoy provable inference scaling laws in the following sense:
Definition 1.
We say that an LLM-based algorithm enjoys a provable inference scaling law for a specific input problem if its success probability (with respect to the inherent randomness of the algorithm) in returning a correct solution to the problem can be boosted arbitrarily close to $100\%$ as its test-time compute grows, provided that certain clearly identified assumptions about the problem and the LLM(s) being used are satisfied.222 For clarity, we emphasize that this definition is stated on a per-instance basis and tailored to an individual input problem, e.g., a multiple-choice question from a question-answering benchmark, or a user query saying “please write a Python function for the quicksort algorithm”.
To this end, we first propose a two-stage knockout-style algorithm that enjoys such provable scaling laws, while requiring a black-box LLM and nothing else (e.g., no external verifier or reward model) for a minimalistic implementation. Given an input problem, the proposed algorithm first generates multiple candidate solutions, and then aggregates them for the final output, via a knockout tournament where pairwise comparisons among the candidates are conducted. We prove theoretically that its failure probability in solving a specific problem decays to zero exponentially (Theorem 1) or by a power law (Theorem 2) with respect to the total number of LLM calls, depending on the specific way of scaling. These guarantees rely on two natural (and arguably minimal) assumptions: (i) the LLM can generate a correct solution with non-zero probability, and (ii) the LLM can do better than a random guess in choosing the right winner between a pair of correct and incorrect solutions.
We further propose a two-stage league-style algorithm that also enjoys a provable scaling law. Unlike the knockout-style algorithm that eliminates a candidate upon loss to a single opponent, the league-style algorithm evaluates each candidate by its average win rate against multiple opponents. We prove theoretically that its failure probability in solving a specific problem decays to zero exponentially (Theorem 3) as its test-time compute scales up, under technical assumptions that are analogous to but more robust than those required by the knockout-style algorithm.
Our practical implementations of both proposed algorithms are efficient and scalable, with support for parallel and distributed computation. Empirical results with the challenging GPQA [30] and MMLU-Pro [37] benchmarks validate our theories, and demonstrate the outstanding scaling properties of both algorithms.
2 A two-stage knockout-style algorithm
This section studies the following two-stage knockout-style algorithm for solving an input problem:
  1. 1.
    Generation. We first generate $N$ candidate solutions, which can run in parallel. In situations where the final answer to the problem contains only a few tokens (e.g., for multiple-choice problems or math calculation), we require that each solution contains a reasoning process or explanation of the rationale that leads to its final answer, which can be elicited by chain-of-thought prompting [38, 17] for example; such information can be useful for enhancing pairwise comparisons in the next stage.
  2. 2.
    Aggregation. We aggregate the candidate solutions via a knockout tournament. At each round of the tournament, the candidates are grouped into pairs randomly, and each pair of candidates are compared for $K$ times. The winner of each pair is the one that is favored for more than $K/2$ times; ties are broken randomly. Only the winners will move on to the next round. The final-round winner at the end of this tournament will be the final output of the algorithm.
A visualization of this process can be found in Figure 1. In a minimalistic implementation, both generation and pairwise comparison can be executed with a single black-box LLM (or an ensemble of multiple LLMs).
Figure 1: A visualization of the proposed two-stage knockout-style algorithm (with $N=8$ in this example).
We next introduce some formal notations, followed by our analysis for both success probability and computational efficiency [4, 15] of the proposed algorithm, to be presented in Sections 2.1 and 2.2 respectively. For concreteness and clarity, our analysis is tailored to the aforementioned minimalistic implementation, but it can certainly be extended to more general cases.
Formal notations.
We let $\mathcal{M}_{\text{gen}}$ and $\mathcal{M}_{\text{comp}}$ denote the probability distribution of the output of one LLM call for generating a candidate solution and for comparing a pair of solutions respectively. Given an input problem $x$, the proposed algorithm first samples $N$ independent candidate solutions
$y_{1},\dots,y_{N}\sim\mathcal{M}_{\text{gen}}(x)$ (1)
during the generation stage. Then, for each pair of candidates $(y,y^{\prime})$ encountered in the knockout stage, the algorithm samples $K$ independent comparison results
$r_{1},\dots,r_{K}\sim\mathcal{M}_{\text{comp}}(x,y,y^{\prime}),$ (2)
and identifies the candidate that is favored by the majority of $\{r_{i}\}_{i\in[K]}$ as the winner. Sampling from $\mathcal{M}_{\text{gen}}$ and $\mathcal{M}_{\text{comp}}$ throughout the algorithm is the sole source of randomness in the following analysis of success probability. The randomness within $\mathcal{M}_{\text{gen}}$ and $\mathcal{M}_{\text{comp}}$ can originate from LLM decoding with a non-zero temperature, the randomized choice of prompting method or LLM backend for each LLM call, among others.
2.1 Analysis of success probability
Our theoretical guarantees for the proposed algorithm rely on the following assumption about the input problem under consideration and the LLM(s) being used.
Assumption 1.
For the input problem $x$, there exists $p_{\text{gen}}>0$ such that
${\mathbb{P}}_{y\sim\mathcal{M}_{\text{gen}}(x)}(y\text{ is a correct solution})\geq p_{\text{gen}}>0.$ (3)
In addition, there exists $p_{\text{comp}}>0.5$ such that, for an arbitrary pair of candidate solutions $(y,y^{\prime})$ where one of them is correct and the other is incorrect, it holds that
${\mathbb{P}}_{r\sim\mathcal{M}_{\text{comp}}(x,y,y^{\prime})}(r\text{ identifies the right winner})\geq p_{\text{comp}}>0.5.$ (4)
In other words, we assume that the LLM can generate a correct solution with non-zero probability, and do better than a random guess in comparing a pair of correct and incorrect solutions. Note that $p_{\text{gen}}$ and $p_{\text{comp}}$ are defined with respect to a specific input problem, not for a distribution of problems or a benchmark.
Scaling up both $N$ and $K$.
Intuitively, as $N$ (the number of initial candidate solutions) and $K$ (the number of times that each pair of solutions involved in the knockout stage are compared) grow, it becomes more likely that (i) there exist initial candidate solutions that are correct ones, and (ii) they tend to be selected as the winners in pairwise comparisons against incorrect solutions, which together lead to the correctness of the final output of the algorithm. This is formalized in the following theorem.
Theorem 1.
If Assumption 1 holds for the input problem, then the failure probability of the proposed two-stage knockout-style algorithm with hyperparameters $N$ and $K$ is bounded by
(final output is ${\mathbb{P}}(\text{final output is incorrect})\leq(1-p_{\text{gen}})^{N}+\lceil\log_{2}N\rceil e^{-2K(p_{\text{comp}}-0.5)^{2}}.$
A proof can be found in Appendix A.1. Notably, the failure probability decays to zero exponentially with respect to the hyperparameters $N$ and $K$ that determine the amount of test-time compute. Another way to interpret this theorem is as follows: for a targeted success probability $1-\delta$ (which can be arbitrarily close to 1 as the failure probability $\delta>0$ approaches zero), it suffices to let $(1-p_{\text{gen}})^{N}\leq\exp(-p_{\text{gen}}N)\leq\delta/2$ and $\lceil\log_{2}N\rceil\exp(-2K(p_{\text{comp}}-0.5)^{2})\leq\delta/2$, namely
$N\geq\frac{1}{p_{\text{gen}}}\log\Big{(}\frac{2}{\delta}\Big{)}\quad\text{and}\quad K\geq\frac{1}{2(p_{\text{comp}}-0.5)^{2}}\log\Big{(}\frac{2\lceil\log_{2}N\rceil}{\delta}\Big{)}.$ (5)
In other words, the hyperparameters $N$ and $K$ have logarithmic dependence on $1/\delta$, and linear dependence on $1/p_{\text{gen}}$ and $1/(p_{\text{comp}}-0.5)^{2}$ respectively.
Scaling up $N$ while $K$ is fixed.
Although Theorem 1 guarantees that an arbitrarily small failure probability can be achieved, it requires the hyperparameter $K$ to be sufficiently large, depending on the value of $p_{\text{comp}}$ that might be unknown a priori in practice. To resolve this concern, we provide an alternative theorem below, which suggests that scaling up $N$ alone is sufficient, even when $K$ is a fixed constant and thus there is still a good chance that the wrong winner is identified when comparing a specific pair of candidates.
To streamline the statement of our theorem, we introduce the notations $\mathcal{M}_{\text{comp},K}$ and $p_{\text{comp},K}$, which generalize $\mathcal{M}_{\text{comp}}$ and $p_{\text{comp}}$ that appear in Eq. (4). Let $\mathcal{M}_{\text{comp},K}$ denote the probability distribution of the comparison result obtained with $K$ independent LLM calls followed by majority voting (with ties broken randomly). Then we have
${\mathbb{P}}_{r\sim\mathcal{M}_{\text{comp},K}(x,y,y^{\prime})}(r\text{ identifies the right winner})\geq p_{\text{comp},K}>0.5,\quad\text{where}$
$p_{\text{comp},K}=\begin{cases}p_{\text{comp}}\quad&\text{if }K=1,\\ \sum_{\ell=\lceil K/2\rceil}^{K}{K\choose\ell}p_{\text{comp}}^{\ell}(1-p_{\text{comp}})^{K-\ell}\quad&\text{if }K\text{ is odd},\\ \sum_{\ell=K/2+1}^{K}{K\choose\ell}p_{\text{comp}}^{\ell}(1-p_{\text{comp}})^{K-\ell}+\frac{1}{2}{K\choose K/2}p_{\text{comp}}^{K/2}(1-p_{\text{comp}})^{K/2}\quad&\text{if }K\text{ is even}.\end{cases}$
Theorem 2.
Suppose that Assumption 1 holds and $N$ is a power of 2. Let $p_{i}$ be the probability that a candidate solution at the $i$-th level of the knockout tournament is correct, where $i=0,1,\dots,\log_{2}N$333 In our notations, the zeroth level of the tournament contains the $N$ initial candidates, the first level contains the $N/2$ winners after the first round of pairwise comparisons, and so on. All candidates within the same level of the knockout tournament have the same probability of being a correct one, due to their symmetric roles. Then
$p_{0}=p_{\text{gen}},\quad\text{and}\quad p_{i+1}\geq p_{i}+(2p_{\text{comp},K}-1)(p_{i}-p_{i}^{2})\quad\text{for}\quad i=0,1,\dots,\log_{2}N-1.$
Consequently, the success probability of the overall algorithm, namely $p_{\log_{2}N}$, converges to 1 as $N$ grows; for any $0.50<\delta<$ 0.5, one has $p_{\log_{2}N}\geq 1-\delta$ as long as
$\log_{2}N\geq\frac{\log\big{(}\max\{\frac{1}{2p_{\text{gen}}},1\}\big{)}}{\log\big{(}1+(p_{\text{comp},K}-0.5)\big{)}}+\frac{\log\big{(}\frac{1}{2\delta}\big{)}}{-\log\big{(}{1-(p_{\text{comp},K}-0.5)}\big{)}}.$
A proof can be found in Appendix A.2. Notably, the linear relationship between $\log_{2}(N)$ and $\log(1/\delta)$ reveals a power-law relationship between the failure probability $\delta$ and the number of candidates $N$.
2.2 Analysis of efficiency
The minimalistic implementation of the proposed knockout-style algorithm starts by generating $N$ candidate solutions with $N$ LLM calls that can run in parallel. Since the number of candidates is reduced by half at each round of the knockout tournament, there is at most $\lceil\log_{2}N\rceil$ rounds in total. For notational convenience, let us assume that $N$ is a power of 2 for the rest of this analysis. At the $i$-th round, there are $N/2^{i}$ pairs of candidates, and each pair need $K$ comparisons; thus a total of $K\times N/2^{i}$ LLM calls are needed, which again can be parallelized.
In sum, the total number of LLM calls required by the two-stage algorithm is
$N+K\times\Big{(}\frac{N}{2}+\frac{N}{2^{2}}+\frac{N}{2^{3}}+\dots+2+1\Big{)}\leq N+K\times N=(K+1)\times N,$
whereas the end-to-end latency, if sufficiently many machines are available, is merely
$T_{\text{gen}}+\log_{2}(N)\times T_{\text{comp}},$
where $T_{\text{gen}}$ and $T_{\text{comp}}$ represent the latency of one LLM call for generating a candidate solution and for comparing a pair of solutions, respectively.
3 A two-stage league-style algorithm
This section introduces a two-stage league-style algorithm that also enjoys a provable scaling law, but under technical assumptions that are different from those required by the knockout-style algorithm.
Motivation: relax Assumption 1.
Recall that Assumption 1 requires that $p_{\text{comp}}>0.5$ holds true for any pair of correct and incorrect solutions, which renders it somewhat restricted and non-robust. This limitation motivates us to propose a league-style algorithm that, as we will show very soon, enjoys a provable scaling law under certain assumptions about the average win rate of a correct candidate solution against a distribution of opponents, rather than against a single one.
The proposed algorithm.
To begin with, we generate $N$ candidate solutions $y_{1},\dots,y_{N}\sim\mathcal{M}_{\text{gen}}(x)$ as before. Then, for each candidate with index $i\in[N]$, we randomly sample $K$ opponents with indices $o_{i}(1),\dots,o_{i}(K)\in[N]\backslash\{i\}$ uniformly and with replacement, conduct one independent pairwise comparison against each opponent, and obtain the responses $r_{i}(j)\sim\mathcal{M}_{\text{comp}}(x,y_{i},y_{o_{i}(j)}),j\in[K]$. The average win rate of each candidate $y_{i}$ is estimated by
$\widehat{\mu}_{i}\coloneqq\frac{1}{K}\sum_{j\in[K]}\phi\big{(}r_{i}(j),y_{i},y_{o_{i}(j)}\big{)},$
where $\phi(r_{i}(j),y_{i},y_{o_{i}(j)})$ denotes the score assigned, based on the response $r_{i}(j)$, to the candidate $y_{i}$ in its comparison against the opponent $y_{o_{i}(j)}$, e.g., 1 for a win, 0 for a loss, and 0.5 for a tie. Finally, the candidate with the highest average win rate $\widehat{\mu}_{i}$ is chosen (with ties broken randomly) as the output of the algorithm.
Regarding computational efficiency, the proposed algorithm (with hyperparameters $N$ and $K$) requires $N$ fully parallelizable LLM calls for the generation stage, and $N\times K$ fully parallelizable LLM calls for the aggregation stage.
Analysis of success probability.
For a solution $y$, we denote its average win rate against $\mathcal{M}_{\text{gen}}$ by
$\mu_{y}\coloneqq{\mathbb{E}}_{y^{\prime}\sim\mathcal{M}_{\text{gen}}(x)}{\mathbb{E}}_{r\sim\mathcal{M}_{\text{comp}}(x,y,y^{\prime})}\phi(r,y,y^{\prime}).$
With this notation in place, we present our key assumption below.
Assumption 2.
For the input problem $x$, there exist $p_{\text{gen}}>0$, $\Delta>0$, and a way of dividing the set $\mathcal{Y}$ of all possible solutions into three disjoint subsets $\mathcal{Y}=\mathcal{Y}_{\text{cs}}\cup\mathcal{Y}_{\text{cw}}\cup\mathcal{Y}_{\text{inc}}$ (where “cs”, “cw” and “inc” in the subscripts stand for “correct-and-strong”, “correct-but-weak” and “incorrect”, respectively), such that
${\mathbb{P}}_{y\sim\mathcal{M}_{\text{gen}}(x)}(y\in\mathcal{Y}_{\text{cs}})\geq p_{\text{gen}}>0\quad\text{and}\quad\min_{y\in\mathcal{Y}_{\text{cs}}}\mu_{y}-\max_{y\in\mathcal{Y}_{\text{inc}}}\mu_{y}\geq\Delta>0.$
In other words, we assume that the LLM can generate a correct-and-strong candidate solution with non-zero probability, whose average win rate against $\mathcal{M}_{\text{gen}}$ is higher (by a strictly positive margin) than that of any incorrect solution. Note that Assumption 2 can be tolerant of systematic errors made by LLMs in comparing certain pairs of candidates, i.e., it may still hold true when there exist a correct solution $y$ and incorrect solution $y^{\prime}$ such that ${\mathbb{E}}_{r\sim\mathcal{M}_{\text{comp}}(x,y,y^{\prime})}\phi(r,y,y^{\prime})<0.5$, whereas Assumption 1 fails in such cases.
Intuitively, if Assumption 2 holds true and the hyperparameters $N$ and $K$ are sufficiently large, then with high probability, (i) there exist initial candidates that are correct-and-strong solutions, and (ii) $\widehat{\mu}_{i}$ is an accurate estimate of $\mu_{y_{i}}$ for each $i\in[N]$. These conditions together lead to a correct final output of the algorithm. We formalize this intuition in the following theorem.
Theorem 3.
If Assumption 2 holds for the input problem, then the failure probability of the league-style algorithm with hyperparameters $N$ and $K$ is bounded by
(final output is ${\mathbb{P}}(\text{final output is incorrect})\leq(1-p_{\text{gen}})^{N}+2Ne^{-{K\Delta^{2}}/{8}}+2Ne^{-{(N-1)\Delta^{2}}/{8}}.$
This theorem, whose proof can be found in Appendix A.3, ensures that the failure probability of the league-style algorithm decays to zero exponentially with respect to $N$ and $K$. Another way to interpret this theorem is as follows: to guarantee success probability $1-\delta$, it suffices to have
$N\geq\max\bigg{\{}\frac{1}{p_{\text{gen}}}\log\Big{(}\frac{3}{\delta}\Big{)},\ \frac{8}{\Delta^{2}}\log\Big{(}\frac{6N}{\delta}\Big{)}+1\bigg{\}}\quad\text{and}\quad K\geq\frac{8}{\Delta^{2}}\log\Big{(}\frac{6N}{\delta}\Big{)}.$
That is, $N$ and $K$ have logarithmic dependence on $1/\delta$, and linear dependence on $\max\{1/p_{\text{gen}},1/\Delta^{2}\}$ and $1/\Delta^{2}$ (up to logarithmic factors) respectively.
Remark 1.
The provable success of the knockout-style algorithm relies on Assumption 1, while that of the league-style algorithm relies on Assumption 2. Although the latter is conceptually more robust than the former, we note that neither assumption is strictly weaker than the other (and thus both algorithms have their unique values). In other words, there exist scenarios where Assumption 1 holds true while Assumption 2 does not, and also scenarios where the reverse is true. Interested readers may refer to Appendix B for some minimal examples.
4 Experiments
We conduct empirical studies in this section to validate the efficacy of the proposed algorithms and confirm the gains from scaling up their test-time compute, while bridging their practical performance with the theories developed in previous sections. Supplementary empirical results can be found in Appendix C.
Datasets.
We use two challenging datasets for our experiments, namely GPQA [30] and MMLU-Pro [37]. GPQA consists of graduate-level multiple-choice questions that pose a challenge even to human experts with PhD degrees in the corresponding domains. It contains over 1000 questions splitted into three categories (“main”, “diamond” and “extended”), all of which are used in our experiments. MMLU-Pro contains 14 categories of multiple-choice questions, some of which require advanced reasoning for obtaining the correct answers. Due to limited computational resources, we use a randomly sampled subset of 100 questions for each category of MMLU-Pro in our experiments, which leads to a total of 1400 questions; we refer to this subset as MMLU-Pro-S in the rest of this work.
Implementations.
We use two open LLMs in our experiments, namely Llama3.1-70B-Instruct (or Llama3.1 for short) [23] and Qwen2.5-72B-Instruct (or Qwen2.5 for short) [28]. We also consider a Mixed option that uses a mixture of both LLMs [34, 48, 12]: during the generation stage, half of the initial candidates are sampled by Llama3.1 and the other half by Qwen2.5; similarly, when a pair of candidates are compared for multiple times during the aggregation stage, half of them are done by Llama3.1 and the other half by Qwen2.5. The rationale is that the capabilities of different LLMs can be complementary to some extent, and thus using a mixture of them can make it more likely that Assumptions 1 and 2 hold true444 To formalize this intuition, consider a minimal scenario with two LLMs denoted by $M_{1}$ and $M_{2}$, and two problems denoted by $x_{1}$ and $x_{2}$. Suppose that $M_{1}$ is good at solving the first problem $x_{1}$ (with $p_{\text{gen}}=0.2$ and $p_{\text{comp}}=0.7$) but incapable of solving $x_{2}$ (with $p_{\text{gen}}=0$ and $p_{\text{comp}}=0.5$), while the reverse is true for $M_{2}$. The magic of using a mixture of two LLMs is that, both problems now satisfy Assumption 1 with $p_{\text{gen}}=(0+0.2)/2=0.1>0$ and $p_{\text{comp}}=(0.5+0.7)/2=0.6>0.5$ (and thus can be solved by our algorithm), whereas only one problem satisfies the assumption when either LLM is used alone. .
Throughout our experiments, the temperature for LLM decoding is set to 0.5 for the generation stage, and 0.1 for pairwise comparisons during the aggregation stage. Our early exploration suggests that these choices strike a good balance between diversity and preciseness in LLM decoding. During the generation stage, we ask the LLM (via zero-shot CoT prompting [17]) to generate a reasoning process first and then its final answer, i.e., the chosen option for a multiple-choice question. For each pairwise comparison during the aggregation stage, we also leverage zero-shot CoT prompting and ask the LLM to think step by step before deciding which solution in the pair is more plausible, unless specified otherwise. The complete prompt templates for both stages can be found in Tables C.3 and C.3 in Appendix C. To account for the positional bias of LLMs [51, 35], we ensure that when a pair of candidates are compared for multiple times, they are placed in one order within the prompt for half of the comparisons, and in the opposite order for the other half.
Unless specified otherwise, for the knockout-style algorithm, we fix the number of comparisons between each pair of candidates encountered in the knockout tournament at $K=4$; for the league-style algorithm, we consider a round-robin (or “all-play-all” [41]) version of it, where 4 pairwise comparisons are conducted between each of (${N\choose 2}$ pairs of initial candidates. To make the proposed algorithms efficient and scalable in practice, we implement them based on AgentScope [9], a multi-agent framework that supports parallel and distributed computation.
In our experiments, we consider a solution to a multiple-choice question as a correct one if its final answer matches the ground-truth answer, and use accuracy as the performance metric for running a (deterministic or randomized) algorithm once on a dataset, i.e., the proportion of problems that the algorithm solves correctly. This metric is, in expectation, equivalent to the mean success probability of the algorithm on the dataset.
4.1 Results for the knockout-style algorithm
Efficacy and scaling properties.
Figure 2: Accuracy versus the number of initial candidates $N$ for the knockout-style algorithm.
In our experiments for the knockout-style algorithm, we scale the number of initial candidates $N$ from 1 to 64. Figure 2 confirms that the accuracy achieved by the knockout-style algorithm improves with $N$ for both datasets and all of Llama3.1, Qwen2.5 and Mixed options. For example, accuracy with the Mixed option shows a $10\%$ improvement on GPQA (from $45\%$ to $55\%$), and a $4.5\%$ improvement on MMLU-Pro-S (from $72\%$ to $76.5\%$). We will see later that improvement in accuracy can be much more significant for certain subsets of either dataset. It is also notable that the Mixed option consistently outperforms Llama3.1 and Qwen2.5, which confirms the previously explained rationales for using a mixture of different LLMs.
But the theorems promise 100% accuracy, don’t they?
Figure 3: The distribution of problems from the GPQA (top) or MMLU-Pro-S (bottom) benchmark, characterized by $\widehat{p}_{\text{gen}}$ and $\widehat{p}_{\text{comp}}$ that are estimated with the empirical results for the knockout-style algorithm using the Llama3.1 (left), Qwen2.5 (middle) or Mixed (right) option. Each plot is annotated with the number of problems satisfying the condition $\widehat{p}_{\text{comp}}>0.5$, $\widehat{p}_{\text{comp}}\leq 0.5$, $\widehat{p}_{\text{gen}}=0$ or $\widehat{p}_{\text{gen}}=1$. To the right of each plot is a histogram for $\widehat{p}_{\text{comp}}$. Each problem is represented by a circle if it is solved correctly by the knockout-style algorithm with $N=64$, and by a cross otherwise. We neglect problems with $\widehat{p}_{\text{gen}}=0$ or $1$, i.e., problems for which the initial candidate solutions are all incorrect or all correct, since there is no way of obtaining meaningful estimate of $\widehat{p}_{\text{comp}}$ for such problems.
Figure 4: Accuracy versus the number of initial candidates $N$ for the knockout-style algorithm, on a filtered subset of problems satisfying $10<\widehat{p}_{\text{gen}}<$ 1 and $\widehat{p}_{\text{comp}}>0.5$. The values of accuracy are calculated with new trials of the algorithm, which are statistically independent of the trials used for estimating $\widehat{p}_{\text{gen}}$ and $\widehat{p}_{\text{comp}}$.
The saturation of accuracy in Figure 2 for large values of $N$ is indeed consistent with our theories. Recall that our theorems developed in Section 2.1 guarantee that the knockout-style algorithm can achieve an arbitrarily high success probability for any input problem that satisfies Assumption 1, namely $p_{\text{gen}}>0$ and $p_{\text{comp}}>0.5$. For a problem that does not, it is still possible that the algorithm has a chance of solving it correctly (since Assumption 1 is a sufficient condition for its success and might not be necessary), but there is no formal guarantee. Consequently, for a benchmark or a distribution of input problems, denoted by $\mathcal{D}$, our algorithm is guaranteed to achieve accuracy (i.e., the average of success probabilities) at least ${\mathbb{P}}_{x\sim\mathcal{D}}(x\text{ satisfies the assumption})$ as its test-time compute scales up.
Motivated by this, we conduct further analysis to enhance our understanding of how the knockout-style algorithm achieves its scaling properties empirically. We start by estimating the parameters $p_{\text{gen}}$ and $p_{\text{comp}}$ in Assumption 1. For each problem from the GPQA and MMLU-Pro-S datasets, we define $\widehat{p}_{\text{gen}}$ as the proportion of the $N=64$ initial candidate solutions with the correct final answer, which serves as a good proxy for $p_{\text{gen}}$. To find a proxy for $p_{\text{comp}}$, we define $\widehat{p}_{\text{comp}}$ by picking all LLM calls for comparing a pair of correct and incorrect solutions throughout the knockout tournament, putting higher weights on the comparison results from later rounds of the tournament, and taking the sum of the weights of all comparisons that identify the right winners [^5] The rationale for weighting the comparison results is explained as follows. In the early rounds of the knockout tournament, the comparison result between a correct-but-weak candidate and an incorrect candidate can cause a negative bias in estimating $p_{\text{comp}}$; similarly, a correct candidate might have a very high win rate against an opponent that is not only incorrect but also very weak, which can cause a positive bias in estimating $p_{\text{comp}}$. In contrast, the correct or incorrect candidates that survive the early rounds of the knockout tournament tend to be stronger ones, which make the comparison results among them (in later rounds of the tournament) more reliable and meaningful for the purpose of estimating $p_{\text{comp}}$. .
Figure 3 characterizes the distribution of problems from the GPQA and MMLU-Pro-S datasets, in terms of $\widehat{p}_{\text{gen}}$ (the X-axis) and $\widehat{p}_{\text{comp}}$ (the Y-axis). On the top half of each scatter plot are the problems with $\widehat{p}_{\text{comp}}>0.5$, most of which are solved correctly by the knockout-style algorithm and represented as circles. These include some problems with small $\widehat{p}_{\text{gen}}$, for which the knockout stage successfully identifies a correct candidate even though the initial candidates are mostly incorrect. We further observe that, compared to Llama3.1 and Qwen2.5, the Mixed option achieves $\widehat{p}_{\text{gen}}>0$ and $\widehat{p}_{\text{comp}}>0.5$ for a larger proportion of problems, which explains its superior accuracy shown in Figure 2. Of course, there also exist problems with $\widehat{p}_{\text{gen}}=0$ or $\widehat{p}_{\text{comp}}\leq 0.5$, which prevent us from boosting the accuracy up to $100\%$ on these benchmarks.
To further consolidate this analysis, we pay special attention to the subset of problems that satisfy $10<\widehat{p}_{\text{gen}}<$ 1 and $\widehat{p}_{\text{comp}}>0.5$. These correspond to the conditions stated in Assumption 1, except that those easy problems with $\widehat{p}_{\text{gen}}=1$ are excluded. We run a new, independent trial of the knockout-style algorithm (with Llama3.1/Qwen2.5/Mixed) on this subset. The results illustrated in Figure 4 confirm that significant improvements in accuracy (e.g., a 25% increase for Mixed) for such problems can be achieved by scaling up the test-time compute of the algorithm, as predicted by our theories. The algorithm does not achieve 100% accuracy on this subset though, since $\widehat{p}_{\text{gen}}$ and $\widehat{p}_{\text{comp}}$ are only approximate estimate of the true underlying $p_{\text{gen}}$ and $p_{\text{comp}}$ respectively, and thus there still exist problems in this subset that violate Assumption 1.
Intuitions: when does Assumption 1 hold?
Figure 5: Accuracy versus $N$ for the knockout-style algorithm on two categories of MMLU-Pro-S.
Interestingly, we observe that for MMLU-Pro-S, the scaling properties of the algorithm vary across different categories of the dataset, and also across the options of which LLM backend(s) to use. For instance, Figure 5 shows that the performance scales well for all of Llama3.1, Qwen2.5 and Mixed in the “engineering” category, while the scaling property of Llama3.1 outperforms the other two options in “philosophy”.
Such results suggest that the gains from scaling up the test-time compute of the knockout-style algorithm depend on the characteristics of both problems and LLM backends. For example, for a reasoning-focused task like “engineering”, LLMs can compare the reasoning processes of two candidate solutions side by side, which provides additional information compared to generating or verifying an individual solution, and thus leads to a large value of $p_{\text{comp}}$ and accurate comparison results. In contrast, for a knowledge-heavy task like “philosophy” that mostly requires memorization of relevant knowledge, one would not expect significant gains from pairwise comparison if the LLM simply does not have the right knowledge embedded within its model weights, in which case $p_{\text{comp}}$ might be close to $0.5$ or worse.
Additional results.
More empirical results for the knockout-style algorithm can be found in Appendix C.1, including the impacts of the hyperparameter $K$ and the use of CoT prompting for the aggregation stage, a comparison with majority voting, and results for each individual category of GPQA and MMLU-Pro-S.
4.2 Results for the league-style algorithm
Efficacy and scaling properties.
Figure 6 shows that the accuracy of the league-style algorithm grows with $N$ until saturation, for both datasets and all of Llama3.1, Qwen2.5 and Mixed options. For example, accuracy of Mixed shows a $8\%$ improvement on GPQA (from $45\%$ to $53\%$) as $N$ grows from 1 to 16, and a $4\%$ improvement (from $72\%$ to $76\%$) on MMLU-Pro-S. The Mixed option consistently outperforms Llama3.1 and Qwen2.5, similar to the case for the knockout-style algorithm. Given the same initial candidates, the league-style algorithm achieves higher accuracy than the knockout-style algorithm (dotted lines) does in some cases and lower in other cases, although the differences are minor in general.
Figure 6: Accuracy versus the number of initial candidates $N$ for the league-style algorithm (LG, solid lines), as well as for the knockout-style algorithm (KO, dotted lines) given the same initial candidates.
Connection to Assumption 2.
Figure 7: The distribution of problems from the GPQA (top) or MMLU-Pro-S (bottom) benchmark, characterized by $\widehat{p}_{\text{gen}}$ and $\widehat{\Delta}$ that are estimated with the empirical results for the league-style algorithm using the Llama3.1 (left), Qwen2.5 (middle) or Mixed (right) option. Each plot is annotated with the number of problems satisfying the condition $\widehat{\Delta}>0$, $\widehat{\Delta}\leq 0$, $\widehat{p}_{\text{gen}}=0$ or $\widehat{p}_{\text{gen}}=1$. To the right of each plot is a histogram for $\widehat{\Delta}$. Each problem is represented by a circle if it is solved correctly by the league-style algorithm with $N=16$, and by a cross otherwise. We neglect problems with $\widehat{p}_{\text{gen}}=0$ or $1$, i.e., problems for which the initial candidate solutions are all incorrect or all correct, since there is no way of obtaining meaningful estimate of $\widehat{\Delta}$ for such problems.
Figure 8: Accuracy versus the number of initial candidates $N$ for the league-style algorithm, on a filtered subset of problems satisfying $10<\widehat{p}_{\text{gen}}<$ 1 and $\widehat{\Delta}>0$. The values of accuracy are calculated with new trials of the algorithm, which are statistically independent of the trials used for estimating $\widehat{p}_{\text{gen}}$ and $\widehat{\Delta}$.
We conduct further analysis that draws connections to the theories developed in Section 3 and enhances our understanding of the empirical performance of the league-style algorithm. Let us start by finding proxies for the $p_{\text{gen}}$ and $\Delta$ parameters in Assumption 2 for each problem of both datasets. For the former, we define $\widehat{p}_{\text{gen}}$ again as the fraction of initial candidate solutions with the correct final answer. For the latter, we define $\widehat{\Delta}\coloneqq\widehat{\mu}_{i_{1}}-\widehat{\mu}_{i_{2}}$, where $i_{1}\coloneqq\arg\max_{i\in[N]:\ y_{i}\text{ is correct}}\widehat{\mu}_{i}$ and $i_{2}\coloneqq\arg\max_{i\in[N]:\ y_{i}\text{ is incorrect}}\widehat{\mu}_{i}$ are the indices for the strongest correct candidate and for the strongest incorrect candidate, respectively. Note that by definition, for any problem with $\widehat{p}_{\text{gen}}\notin\{0,1\}$, the league-style algorithm returns a correct solution to the problem if and only if $\widehat{\Delta}>0$.
Figure 7 characterizes the distribution of problems from the GPQA and MMLU-Pro-S datasets, in terms of $\widehat{p}_{\text{gen}}$ and $\widehat{\Delta}$. It is noteworthy that there exists a non-trivial proportion of problems for which $\widehat{p}_{\text{gen}}$ is fairly small (i.e., most of the initial candidates are incorrect), while the league-style aggregation stage manages to attain a positive $\widehat{\Delta}$ and thus identify a correct candidate for the final output, which demonstrates the benefits of the league-style aggregation method. We also pay special attention to the subset of problems satisfying $10<\widehat{p}_{\text{gen}}<$ 1 and $\widehat{\Delta}>0$, which correspond to the conditions stated in Assumption 2, except that those easy problems with $\widehat{p}_{\text{gen}}=1$ are excluded. We run a new, independent trial of the league-style algorithm (with Llama3.1/Qwen2.5/Mixed) on this subset. Figure 8 confirms that accuracy improves with $N$ significantly on this subset (e.g., from 57% to 82% for Mixed), as predicted by our theories.
Efficacy of subsampling opponents.
While all previous experiments consider the round-robin version of the league-style algorithm, we also wonder if it is feasible to improve its computational efficiency by estimating the average win rate of each candidate with $M<N-1$ subsampled opponents, while maintaining its accuracy. The empirical results in Figure 9 provide a positive answer and match what our theories in Section 3 predict: (i) accuracy initially increases with $M$, which confirms the benefits of comparing each candidate with multiple opponents; (ii) once $M$ exceeds a threshold (around 4 or 5) that is much smaller than $N=16$, accuracy saturates around the level achieved by the round-robin version, but at a lower computational cost.
Figure 9: Accuracy versus $M$, the number of subsampled opponents for each candidate, for the league-style algorithm with $N=16$.
5 Related works
There has been a surge of interest in enhancing problem solving with LLMs by various strategies of scaling up test-time compute, including those mentioned earlier in Section 1. Extensive literature review on this topic can be found in some recent surveys [39, 47, 11], and thus we will focus our discussions in this section mainly on prior works that are closely related to ours.
A provable inference scaling law (in the sense of Definition 1) has been developed for majority voting in the literature [3, 42], which essentially requires two assumptions. (i) First, it should be feasible to divide the candidate solutions into several groups and have a meaningful count for each group. This is true for, say, multiple-choice questions, but not for many other tasks where all candidate solutions are distinct, such as open-ended writing. In contrast, both knockout-style and league-style algorithms proposed in our work, where selecting a good solution out of many candidates is reduced to a series of pairwise comparisons (or two-choice questions), can be applied to a wider range of tasks in a unified manner. (ii) Second, majority voting requires that the probability of generating a solution that belongs to the correct group, akin to $p_{\text{gen}}$ in our notations, should be larger than that of any other group; otherwise, it has been proved that the success probability of majority voting will converge to zero instead as its test-time compute grows [3]. In comparison, our proposed algorithms only require $p_{\text{gen}}>0$, though with additional assumption about LLMs’ capabilities of conducting pairwise comparisons.
Deriving a provable scaling law for the Best-of-N approach is straightforward, provided that a perfect verifier is available for picking a correct solution out of $N$ candidates: if one LLM call generates a correct solution with probability at least $p_{\text{gen}}>0$, then the failure probability of Best-of-N is upper bounded by $(1-p_{\text{gen}})^{N}$, which converges to zero as the number of samples $N$ grows. One obvious limitation, of course, is that verifiers are unavailable or imperfect in many practical scenarios, which can hinder the performance of Best-of-N as its test-time compute grows [5, 32].
Our algorithm design in this work has drawn inspiration from various areas. For example, the essential idea of pairwise comparison has been prominent in LLM alignment [1, 27, 29], as well as in the LLM-as-a-judge approach [51, 19] that has been used mostly for the purpose of scalable and automatic LLM evaluation, but also as a key component of test-time strategies in some prior works (cf. Section 5.4 of [19]). Although it is possible to verify, score or refine a solution by itself [13, 24, 7, 10], intuitively it is often much easier (for LLMs or human) to detect the errors or hallucinations in an incorrect solution when it is placed right next to a correct one, or evaluate the quality of a solution by comparing it to another one. The knockout and league tournaments have also been investigated in prior LLM research [16, 22, 50, 18, 12], albeit with purposes or concrete implementations that are different from ours. Given this context, we remark that the main novelty and contributions in this work are perhaps less about the proposed two-stage algorithms themselves, but rather more about developing rigorous and theoretical understanding of their underlying assumptions and efficacy (via clearly identifying sufficient conditions for boosting their success probability up to $100\%$ and formally deriving quantitative bounds for their computational and sample complexities), and demonstrating their promising empirical performance through extensive experiments.
With the ongoing advancements of LLMs (such as OpenAI’s o1 [26, 47]), we anticipate that stronger LLMs in the future will achieve higher values of $p_{\text{gen}}$, $p_{\text{comp}}$ and $\Delta$ in Assumptions 1 and 2 for the same input problem, which will reduce the number of LLM calls required by our algorithms, according to the proposed theorems. Furthermore, these technical assumptions will become valid for more challenging problems that currently violate them (due to limitations of the existing LLMs), which will enable the proposed algorithms to achieve success in broader scenarios.
6 Discussions
We conclude by discussing some limitations of the current work as well as directions for future work.
Solving hard problems by task decomposition.
Our theorems are not directly applicable to challenging scenarios where the LLM simply has zero chance of generating a correct solution to the input problem in one shot, i.e., when $p_{\text{gen}}=0$. For such scenarios, a generic approach adopted by practical LLM-based algorithms and agentic workflows [4, 46, 43] is to decompose the original problem into smaller sub-problems and solve each of them. As long as the sub-problems satisfy the technical assumptions in our theories, we can apply the proposed algorithms to each of them. This approach can offer not only formal guarantees for the success probability of solving the original problem, but also higher efficiency compared to solving it directly.
To see this, consider a scenario where solving the original problem requires solving all $S\geq 1$ sub-problems correctly, and each sub-problem satisfies Assumption 1 with parameters $p_{\text{gen}}>0$ and $p_{\text{comp}}>0.5$. Directly solving all $S$ sub-problems has a exponentially small success probability $p_{\text{gen}}^{S}$, and thus generating a correct solution alone already requires $\Omega((1/p_{\text{gen}})^{S})$ attempts, not to mention identifying which attempt is successful. In contrast, by conducting task decomposition and applying the knockout-style algorithm (with hyperparameters $N$ and $K$) to each sub-problem, an overall success probability of $1-\delta$ for solving the original problem can be guaranteed as long as the failure probability for each sub-problem is bounded by $\delta/S$, thanks to the union bound. According to Eq. (5), this is guaranteed with
$N\geq\frac{1}{p_{\text{gen}}}\log\Big{(}\frac{2S}{\delta}\Big{)}\quad\text{and}\quad K\geq\frac{1}{2(p_{\text{comp}}-0.5)^{2}}\log\Big{(}\frac{2\lceil\log_{2}N\rceil S}{\delta}\Big{)}.$
Note that $N$ and $K$ only have logarithmic dependence on the number of sub-problems $S$. The total number of LLM calls with this approach is $(K+1)\times N\times S$ (cf. Section 2.2), which grows with $S$ linearly, up to logarithmic factors.
More general problem settings.
One limitation of the current work is that, we have focused on the binary metric of correctness throughout, and conducted experiments with only two benchmarks (due to limited computational resources). Future work may extend the theories to more general performance metrics, and test the proposed algorithms in broader scenarios.
Another different problem setting that we have not investigated is about optimal allocation of test-time compute [31, 3, 6, 48]: given a fixed budget of test-time compute and a set of problems to be solved, is there a method for automatically finding the optimal configuration for each problem (e.g., the values of $N$ and $K$ in our algorithms), in order to maximize the overall performance, such as the number of solved problems? Achieving this for the proposed algorithms will likely be a fruitful direction for future research.
Finally, let us conclude with some discussions about anytime algorithms [40]. In many real-world scenarios, the available amount of test-time compute is adaptive and unknown a priori. To address such cases, we can easily convert the knockout-style algorithm to an “anytime” variant that does not require pre-specifying $N$. For example, the algorithm might start with 4 candidate solutions and choose the winner via a knockout tournament. If more test-time compute is allowed (e.g., the user is not eagerly requesting the solution, or more computational resources become available), then the algorithm can launch another tournament with 4 freshly sampled candidates, the winner of which will compete with the winner of the previous tournament. This complete process is indeed equivalent to a single tournament with $N=4+4=8$. Such a process can continue until the user finally requests the solution; the eventual value of $N$ is determined online and automatically achieves the maximum value allowed by the available test-time compute. Similarly, the league-style algorithm can be converted to an anytime variant, where the total number of candidates and/or the number of comparisons for each candidate increase gradually as more test-time compute becomes available. It would be interesting future work to investigate such anytime algorithms from a theoretical perspective, or apply these methods to solving practical problems in online scenarios.
Appendix A Proofs of main theorems
A.1 Proof of Theorem 1
To begin with, we have a straightforward analysis for the failure probability of the generation stage of the algorithm, where $N$ candidate solutions are sampled independently:
(no candidate solution is ${\mathbb{P}}(\text{no candidate solution is correct})\leq(1-p_{\text{gen}})^{N}.$
As for the knockout stage, let us first consider a single pair of correct and incorrect candidate solutions. Recall that they are compared for $K$ times with $K$ LLM calls (followed by majority voting), and each LLM call identifies the correct candidate solution as the winner with probability $\mu\geq p_{\text{comp}}>0.5$ by assumption. Therefore, the failure probability of comparing this pair of candidates can be bounded as follows, where $X_{i}$ denotes an independent Bernoulli random variable with mean $\mu$:
(failure of ${\mathbb{P}}(\text{failure of comparison})$ $\leq{\mathbb{P}}\Big{(}\sum_{i\in[K]}X_{i}\leq\frac{K}{2}\Big{)}={\mathbb{P}}\Big{(}\frac{1}{K}\sum_{i\in[K]}X_{i}\leq 0.5\Big{)}$
$={\mathbb{P}}\Big{(}\frac{1}{K}\sum_{i\in[K]}X_{i}-\mu\leq-(\mu-0.5)\Big{)}$
$\leq\exp\Big{(}-2K(\mu-0.5)^{2}\Big{)}\leq\exp\Big{(}-2K(p_{\text{comp}}-0.5)^{2}\Big{)}.$
Here we use Hoeffding’s inequality [33] in the last line.
Now we are ready to control the failure probability of the complete knockout stage. Let us condition on the event that the generation stage succeeds, i.e., there is at least one initial candidate solution that is correct. We arbitrarily pick a correct candidate, and focus on its path to the final output of the algorithm in the binary tree visualized in Figure 1. We claim that, with high probability, the comparison (with $K$ LLM calls) for each pair along this path yields the correct outcome. This can be proved by induction: for each pair along this path, if one of the input candidates (which is the output of the previous pairwise comparison on this same path) is correct, then the output of comparing this pair will also be correct with a failure probability no greater than $\exp(-2K(p_{\text{comp}}-0.5)^{2})$, regardless of whether the other input candidate is correct or not. By taking a union bound over the failure events along this path with $\lceil\log_{2}N\rceil$ pairs to be compared, we claim that the comparison for each pair along this path yields the correct outcome (which immediately implies that the final output of the algorithm is correct), with a failure probability no greater than $\lceil\log_{2}N\rceil\exp(-2K(p_{\text{comp}}-0.5)^{2})$.
Finally, taking a union bound over the failure events of both stages of the algorithm completes our proof of Theorem 1.
Remark 2.
There exists analysis in the literature of top-k ranking (e.g., Section 4.1 of [25]) that are similar to our analysis for the knockout stage. We choose to present our own version here to make our work more self-contained and complete.
A.2 Proof of Theorem 2
For the first part of the theorem, we can derive $p_{i+1}$ from $p_{i}$ as follows. Notice that a candidate at the $(i+1)$-th level of the knockout tournament is the winner of pairwise comparison between a pair of statistically independent candidates at the $i$-th level. Thus, the winner is a correct solution if both candidates of the pair are correct, or only one of them is correct and happens (with probability at least $p_{\text{comp},K}$ by assumption) to be chosen as the winner. Therefore,
$p_{i+1}$ $\geq p_{i}^{2}+2p_{i}(1-p_{i})p_{\text{comp},K}=p_{i}^{2}+2p_{\text{comp},K}(p_{i}-p_{i}^{2})=p_{i}-p_{i}+p_{i}^{2}+2p_{\text{comp},K}(p_{i}-p_{i}^{2})$
$=p_{i}+(2p_{\text{comp},K}-1)(p_{i}-p_{i}^{2}).$
This implies $p_{i+1}>p_{i}$, as long as $p_{\text{comp},K}>0.5$ and $p_{i}<1$.
For the second part of the theorem, we consider the convergence of $\{p_{i}\}$ in two cases: when it is still below $0.5$, and when it has exceeded $0.5$.
  • If $p_{i}<0.5$, then $0.51-p_{i}>$ 0.5, and
    $p_{i+1}$ $\geq p_{i}+(2p_{\text{comp},K}-1)(1-p_{i})p_{i}$
    $>p_{i}+(p_{\text{comp},K}-0.5)p_{i}$
    $=\Big{(}1+(p_{\text{comp},K}-0.5)\Big{)}p_{i}.$
    In other words, the sequence $\{p_{i}\}$ grows exponentially when it is below $0.5$, and
    $p_{J}\geq p_{\text{gen}}\Big{(}1+(p_{\text{comp},K}-0.5)\Big{)}^{J}\geq 0.5\quad\text{as long as}\quad J\geq\frac{\log\big{(}\max\{\frac{1}{2p_{\text{gen}}},1\}\big{)}}{\log\big{(}1+(p_{\text{comp},K}-0.5)\big{)}}.$ (6)
  • For any $i>J$ and hence $p_{i}\geq 0.5$, we have
    $1-p_{i+1}$ $\leq 1-p_{i}-(2p_{\text{comp},K}-1)p_{i}(1-p_{i})$
    $\leq 1-p_{i}-(p_{\text{comp},K}-0.5)(1-p_{i})$
    $=\Big{(}1-(p_{\text{comp},K}-0.5)\Big{)}(1-p_{i}).$
    In other words, the sequence $\{1-p_{i}\}$ converges to 0, and
    $1-p_{i}\leq(1-p_{J})\Big{(}1-(p_{\text{comp},K}-0.5)\Big{)}^{i-J}\leq\frac{1}{2}\Big{(}1-(p_{\text{comp},K}-0.5)\Big{)}^{i-J}\leq\delta$
    as long $\text{as long as}\quad i-J\geq\frac{\log\big{(}\frac{1}{2\delta}\big{)}}{-\log\big{(}{1-(p_{\text{comp},K}-0.5)}\big{)}}.$ (7)
Putting Eq. (6) and Eq. (7) together concludes our proof of the theorem.
A.3 Proof of Theorem 3
To begin with, we have a straightforward analysis for the generation stage:
${\mathbb{P}}\Big{(}y_{i}\notin\mathcal{Y}_{\text{cs}},\forall i\in[N]\Big{)}\leq(1-p_{\text{gen}})^{N}.$
For the aggregation stage, we aim to show that for each $i\in[N]$, the estimated average win rate $\widehat{\mu}_{i}$ calculated within the algorithm is close to its average win rate against $\mathcal{M}_{\text{gen}}$, denoted by $\mu_{i}\coloneqq\mu_{y_{i}}$. To see this, let us recall the definitions of $\mu_{i}$ and $\widehat{\mu}_{i}$, as well as introduce a new notation $\widetilde{\mu}_{i}$:
$\mu_{i}$ $\coloneqq{\mathbb{E}}_{y^{\prime}\sim\mathcal{M}_{\text{gen}}(x)}{\mathbb{E}}_{r\sim\mathcal{M}_{\text{comp}}(x,y_{i},y^{\prime})}\phi(r,y_{i},y^{\prime}),$
$\widetilde{\mu}_{i}$ $\coloneqq\frac{1}{N-1}\sum_{j\in[N]\backslash\{i\}}{\mathbb{E}}_{r\sim\mathcal{M}_{\text{comp}}(x,y_{i},y_{j})}\phi(r,y_{i},y_{j})$
$={\mathbb{E}}_{y^{\prime}\sim\text{Unif}(y_{j},j\in[N]\backslash\{i\})}{\mathbb{E}}_{r\sim\mathcal{M}_{\text{comp}}(x,y_{i},y^{\prime})}\phi(r,y_{i},y^{\prime}),$
$\widehat{\mu}_{i}$ $\coloneqq\frac{1}{K}\sum_{j\in[K]}\phi(r_{i}(j),y_{i},y_{o_{i}(j)}).$
Note that in the last line, $y_{o_{i}(j)}\sim\text{Unif}(y_{j},j\in[N]\backslash\{i\})$, and $r_{i}(j)\sim\mathcal{M}_{\text{comp}}(x,y_{i},y_{o_{i}(j)})$. By Hoeffding’s inequality, we have the following for a specific $i\in[N]$:
${\mathbb{P}}\Big{(}|\widehat{\mu}_{i}-\widetilde{\mu}_{i}|\geq\frac{\Delta}{4}\Big{)}$ $={\mathbb{P}}\bigg{(}\Big{|}\frac{1}{K}\sum_{j\in[K]}\phi(r_{i}(j),y_{i},y_{o_{i}(j)})-\widetilde{\mu}_{i}\Big{|}\geq\frac{\Delta}{4}\bigg{)}\leq 2\exp\Big{(}-\frac{K\Delta^{2}}{8}\Big{)},$
${\mathbb{P}}\Big{(}|\widetilde{\mu}_{i}-\mu_{i}|\geq\frac{\Delta}{4}\Big{)}$ $={\mathbb{P}}\bigg{(}\Big{|}\frac{1}{N-1}\sum_{j\in[N]\backslash\{i\}}{\mathbb{E}}_{r\sim\mathcal{M}_{\text{comp}}(x,y_{i},y_{j})}\phi(r,y_{i},y_{j})-\mu_{i}\Big{|}\geq\frac{\Delta}{4}\bigg{)}\leq 2\exp\Big{(}-\frac{(N-1)\Delta^{2}}{8}\Big{)}.$
These, together with the fact that $|\widehat{\mu}_{i}-\mu_{i}|\leq|\widehat{\mu}_{i}-\widetilde{\mu}_{i}|+|\widetilde{\mu}_{i}-\mu_{i}|$, implies that
${\mathbb{P}}\Big{(}|\widehat{\mu}_{i}-\mu_{i}|\geq\frac{\Delta}{2}\Big{)}\leq{\mathbb{P}}\Big{(}|\widehat{\mu}_{i}-\widetilde{\mu}_{i}|\geq\frac{\Delta}{4}\Big{)}+{\mathbb{P}}\Big{(}|\widetilde{\mu}_{i}-\mu_{i}|\geq\frac{\Delta}{4}\Big{)}\leq 2\exp\Big{(}-\frac{K\Delta^{2}}{8}\Big{)}+2\exp\Big{(}-\frac{(N-1)\Delta^{2}}{8}\Big{)}.$
Finally, taking a union bound over $i\in[N]$ and over both stages of the league-style algorithm, we have the following: with probability at least
$1-(1-p_{\text{gen}})^{N}-2N\exp\Big{(}-\frac{K\Delta^{2}}{8}\Big{)}-2N\exp\Big{(}-\frac{(N-1)\Delta^{2}}{8}\Big{)},$
there exists some $i\in[N]$ such that $y_{i}\in\mathcal{Y}_{\text{cs}}$, and $|\widehat{\mu}_{j}-\mu_{j}|<\Delta/2$ for all $j\in[N]$. These conditions, together with the assumption that $\min_{y\in\mathcal{Y}_{\text{cs}}}\mu_{y}-\max_{y\in\mathcal{Y}_{\text{inc}}}\mu_{y}\geq\Delta$, guarantee that the final output of the algorithm is a correct solution.
Appendix B Examples for understanding and comparing the assumptions
This section presents some minimal examples for assisting our understanding of Assumptions 1 and 2, and in particular, for comparing the condition $p_{\text{comp}}>0.5$ stated in the former and $\Delta>0$ stated in the latter. For simplicity, we assume that the set of all possible candidate solutions returned by the generation stage, denoted by $\mathcal{Y}$, has a small number of unique elements, e.g., $\mathcal{Y}=\{A,B,C\}$. We use the notation $p_{A}\coloneqq{\mathbb{P}}_{y\sim\mathcal{M}_{\text{gen}}(x)}(y=A)$, and let ${\mathbb{P}}(A\succ B)$ denote the probability that one comparison between $A$ and $B$ identifies the former as the winner. When two identical candidates are compared, we assume that tie is broken randomly and thus either candidate wins with probability $0.5$. All average win rates involved in these examples are calculated with respect to the distribution $\mathcal{M}_{\text{gen}}$.
Example 1.
We demonstrate a scenario where both Assumptions 1 and 2 hold, and there is a correspondence between the parameter $p_{\text{comp}}$ in the former and $\Delta$ in the latter. Suppose that $\mathcal{Y}=\{A,B\}$, where $A$ is correct and $B$ is incorrect. In addition, $p_{A}=\alpha,p_{B}=1-\alpha$, and ${\mathbb{P}}(A\succ B)=p_{\text{comp}}>0.5$. Then we can calculate the average win rate of each candidate as follows:
$\mu_{A}$ $=p_{A}\times 0.5+p_{B}\times p_{\text{comp}}=0.5\times\alpha+p_{\text{comp}}\times(1-\alpha)=(0.5-p_{\text{comp}})\times\alpha+p_{\text{comp}},$
$\mu_{B}$ $=p_{B}\times 0.5+p_{A}\times(1-p_{\text{comp}})=0.5\times(1-\alpha)+(1-p_{\text{comp}})\times\alpha=(0.5-p_{\text{comp}})\times\alpha+0.5,$
which implies
$\Delta=\mu_{A}-\mu_{B}=p_{\text{comp}}-0.5>0$
regardless of the value of $\alpha$.
Example 2.
We demonstrate a scenario where both Assumptions 1 and 2 hold, but the parameter $\Delta$ in the latter can be much smaller than $p_{\text{comp}}-0.5$ in the former. Suppose that $\mathcal{Y}=\{A,B,C\}$, where only $A$ is correct. In addition, $p_{A}=p_{B}=\alpha$, $p_{C}=1-2\alpha$, ${\mathbb{P}}(A\succ B)=p_{\text{comp}}>0.5$, and ${\mathbb{P}}(A\succ C)={\mathbb{P}}(B\succ C)=0.9$. Then we can calculate the average win rate of each candidate as follows:
$\mu_{A}$ $=p_{A}\times 0.5+p_{B}\times p_{\text{comp}}+p_{C}\times 0.9$ $=\alpha\times(0.5+p_{\text{comp}})+(1-2\alpha)\times 0.9,$
$\mu_{B}$ $=p_{A}\times(1-p_{\text{comp}})+p_{B}\times 0.5+p_{C}\times 0.9$ $=\alpha\times(1.5-p_{\text{comp}})+(1-2\alpha)\times 0.9,$
$\mu_{C}$ $=p_{A}\times 0.1+p_{B}\times 0.1+p_{C}\times 0.5$ $=\alpha\times 0.2+(1-2\alpha)\times 0.5,$
which implies
$\Delta=\mu_{A}-\mu_{B}=2\alpha\times(p_{\text{comp}}-0.5).$
As a result, $\Delta>0$ can be much smaller than $p_{\text{comp}}-0.5$ if $\alpha$ is small.
Example 3.
We demonstrate a scenario where Assumption 1 holds true but Assumption 2 does not. Suppose that $\mathcal{Y}=\{A,B,C\}$, where only $A$ is correct. In addition, $p_{A}=0.2,p_{B}=0.2,p_{C}=0.6$, ${\mathbb{P}}(A\succ B)={\mathbb{P}}(A\succ C)=0.6$, and ${\mathbb{P}}(B\succ C)=0.9$, which satisfies Assumption 1. Then we have
$\mu_{A}$ $=p_{A}\times 0.5+p_{B}\times 0.6+p_{C}\times 0.6=0.58,$
$\mu_{B}$ $=p_{A}\times 0.4+p_{B}\times 0.5+p_{C}\times 0.9=0.72,$
$\mu_{C}$ $=p_{A}\times 0.4+p_{B}\times 0.1+p_{C}\times 0.5=0.40.$
In other words, the average win rate of the only correct solution $A$ is lower than that of an incorrect solution $B$, which violates Assumption 2.
Example 4.
We demonstrate a scenario where Assumption 2 holds true but Assumption 1 does not. Suppose that $\mathcal{Y}=\{A,B,C\}$, where only $A$ is correct. In addition, $p_{A}=0.2,p_{B}=0.2,p_{C}=0.6$, ${\mathbb{P}}(A\succ B)=0.4,{\mathbb{P}}(A\succ C)=0.9$, and ${\mathbb{P}}(B\succ C)=0.5$, which violates Assumption 1. However, we have
$\mu_{A}$ $=p_{A}\times 0.5+p_{B}\times 0.4+p_{C}\times 0.9=0.72,$
$\mu_{B}$ $=p_{A}\times 0.6+p_{B}\times 0.5+p_{C}\times 0.5=0.52,$
$\mu_{C}$ $=p_{A}\times 0.1+p_{B}\times 0.5+p_{C}\times 0.5=0.42,$
which satisfies Assumption 2 since the only correct solution $A$ has the highest average win rate.
Appendix C Additional empirical results
C.1 Results for the knockout-style algorithm
Ablation: the impact of $K$.
The results in Figure 10 suggest that the performance of the knockout-style algorithm is insensitive to $K$ (the number of times that each pair of candidates are compared) in the setting of our experiments, as long as $K\geq 2$ for Llama3.1 and Qwen2.5, or $K\geq 4$ for Mixed. This is mainly due to our choice of a small temperature (0.1) for LLM calls that conduct pairwise comparisons. For Llama3.1 and Qwen2.5, $K=2$ suffices to cover all prompting options, i.e., the order in which two candidate solutions are placed within the prompt. Similarly, for Mixed, $K=4$ suffices to cover both prompting options and both LLM backends.
Figure 10: The impact of $K$ for the knockout-style algorithm.
Ablation: the impact of CoT prompting for pairwise comparison.
Figure 11 confirms the benefits of using zero-shot chain-of-thought prompting for the aggregation stage of the knockout-style algorithm (versus prompting the LLM to answer directly which solution is preferred), especially as the test-time compute scales up. This matches the intuition that CoT prompting improves LLMs’ performance in conducting pairwise comparisons.
Figure 11: The advantages of zero-shot chain-of-thought prompting for pairwise comparisons, versus prompting the LLM to answer directly which solution is preferred (dashed lines), during the aggregation stage of the knockout-style algorithm.
A comparison with majority voting.
Figure 12 shows that the knockout-style algorithm achieve higher accuracy than majority voting does in some cases, e.g., for the Qwen2.5 and Mixed options in GPQA, but not in all cases. Note that the knockout-style algorithm requires more test-time compute than majority voting does in this comparison; nonetheless, based on the trends shown in the figure, it is most likely that majority voting will not benefit much from further increasing its budget [3].
Figure 12: A comparison between the knockout-style algorithm (solid) and majority voting (dotted).
Results for each category of GPQA and MMLU-Pro-S.
Figure 13 includes empirical results of the knockout-style algorithm for each category of GPQA, while Figures 14 and 15 include those for MMLU-Pro-S.
Figure 13: Empirical results of the knockout-style algorithm for each category of GPQA.
Figure 14: Empirical results of the knockout-style algorithm for each category of MMLU-Pro-S (Part 1).
Figure 15: Empirical results of the knockout-style algorithm for each category of MMLU-Pro-S (Part 2).
C.2 Results for the league-style algorithm
A closer look at both algorithms and their differences.
Figure 16 provides a detailed comparison between the empirical performance of both algorithms. It characterizes the distribution of GPQA and MMLU-Pro-S problems in terms of $\widehat{p}_{\text{comp}}$ from the knockout-style algorithm and $\widehat{\Delta}$ from the league-style algorithm, and provides the concrete number of problems that one algorithm solves correctly/incorrectly and the other algorithm solves correctly/incorrectly.
Results for each category of GPQA and MMLU-Pro-S.
Figure 17 includes empirical results of the league-style algorithm for each category of GPQA, while Figures 18 and 19 include those for MMLU-Pro-S.
Figure 16: The distribution of problems from the GPQA (top) or MMLU-Pro-S (bottom) benchmark, characterized by $\widehat{p}_{\text{comp}}$ from the knockout-style algorithm and $\widehat{\Delta}$ from the league-style algorithm (both with $N=16$) using the Llama3.1 (left), Qwen2.5 (middle) or Mixed (right) option. The following abbreviations are used for the legend: K — knockout, L — league, R — right, W — wrong. For example, “KW-LR (# = 66)” means that there are 66 problems for which the knockout-style algorithm did wrong while the league-style algorithm did right.
Figure 17: Empirical results of the league-style algorithm for each category of GPQA.
Figure 18: Empirical results of the league-style algorithm for each category of MMLU-Pro-S (Part 1).
Figure 19: Empirical results of the league-style algorithm for each category of MMLU-Pro-S (Part 2).
C.3 Prompt templates
Tables C.3 and C.3 include the prompt templates used in our experiments for generating a candidate solution and for comparing a pair of candidates, respectively. Some parts of our prompts, as well as code for parsing LLMs’ responses and extracting the answers for evaluation, are modified from those in the official GitHub repository of MMLU-Pro [^6] https://github.com/TIGER-AI-Lab/MMLU-Pro/tree/main.
Table 1: The adopted prompt template for generating a candidate solution.
% System prompt Please read the following multiple-choice questions and provide the most likely correct answer based on the options given. % User prompt # Question {question} # Output Format “‘ <reason>your step-by-step reasoning proecss</reason> <answer>"the answer is (X)" where X is the correct letter choice</answer> “‘
Table 2: The adopted prompt template for pairwise comparison.
% System prompt You are an impartial Judge. Given a question and two candidate solutions, your task is to choose which solution answer the question better. Your judgment should be unbiased, without favoring either Solution 1 or 2. % User prompt —- QUESTION —- {question} —- Solution 1 —- {candidate_a} —- Solution 2 —- {candidate_b} —- OUTPUT FORMAT —- “‘ <compare>compare both candidate solutions step-by-step thoroughly, and double check if there are mistakes in either solution</compare> <winner>Solution 1 or Solution 2 or Tie</winner> “‘